# Do LLMs Have Distinct and Consistent Personality? TRAIT: Personality Testset designed for LLMs with Psychometrics

### v1

@@ 3.3 Auditing TRAIT
We test our TRAIT’s quality with two psychological professionals, asking to guess the binary level (High or Low) of option paired with situation and query (random baseline gives an accuracy of 50%). Due to the cost limit, we subsample 200 items for human validation, and the accuracy is 97.5% confirming the quality of the data. More details are in Appendix I.
@@
To confirm that TRAIT is more valid and reliable in assessing personality of LLM than existing baselines, we test all the validity and the reliability introduced in §2.3 on TRAIT. As shown in Table 1, TRAIT achieves the highest marks in both validity and reliability among the personality tests. Specifically, it significantly outperforms all baselines with a 0.3% of refusal rate and shows a clear improvement in average sensitivity, showing improvements of more than 9.3% compared to self-assessments.
@@
To further test the fidelity of TRAIT, we fine-tune a multi-task classification model with TRAIT. T-evaluator can do two tasks differentiated by the instruction: 1) Trait classification: identify the most relevant personality trait from the given text (8 traits), and 2) Level classification: determine the level of given trait revealed in given input (High or Low, 2 classes). We use a concatenation of situation, question, and one of the options as a given sentence and train classifier to generate categorized trait (e.g., Extraversion) or the level (e.g., high). For more training details, see Appendix D.1.
In TRAIT, each personality description is augmented to five different situations, enabling the observation of variations in the models’ responses according to the context. We report the number of high and low personality responses selected by the model when it is presented with the same query across five different scenarios, based on the responses from eight models. Table 5 shows that the models often select two or three high personality responses, not selecting zero or five responses when given the situation. This implies that model personality highly relies on the situation, which is intuitive — humans also change their behavior based on the context they are in Sauerberger and Funder (2017). Specifically, in Agreeableness and Extraversion, the models often choose two or three high personality responses, which are more than 50%. Conversely, for Narcissism, the models commonly choose zero or five high personality responses (61.2%). To see more qualitative results, see Appendix J.
@@ 3.4 Diverse and Detailed Scenarios are Needed when Measuring LLM Personality
Table 5: (#high, #low) indicates the number of high and low responses in five questions rooted in the same persona description but featuring different scenarios, and the other columns indicate the ratio (%) of each cases per the trait.
@@ 4.1 Do LLMs have Distinct Personality?
We first test the personality scores of the nine highly capable models — GPT-4, Claude-sonnet, GPT-3.5, Mistral-7B, Mistral-7B-inst, Llama2-7B, Llama3-8B, Llama3-8B-inst and gemma-2B. Figure 4 shows the scores of the models on eight personality traits. In general, we observe that aligned LLMs — GPT-4, Claude-sonnet, GPT-3.5, Mistral-7B-inst, and Llama3-8B-inst — shows higher scores in agreeableness (78.3 vs 66.7) and conscientiousness (91.0 vs 81.7) and lower scores in openness (56.3 vs 67.8) and extraversion (32.8 vs 46.9). Given that these aligned models are fine-tuned to act as an assistant, such tendency is interesting as the existing study on human subjects claims that the group of teaching assistants in school exhibit higher Agreeableness and Conscientiousness than the average people, while lower Openness and Neuroticism Dočkalová et al. (2023). The aligned models also show lower scores in the Dark Triad compared to the other four models (9.3 vs 27.0), especially compared to Llama2-7B which shows high scores of 42-48 on these traits. Alignment tuning generally targets to reduce the harmfulness of LLMs, and we speculate this objective leads to low scores in the Dark Triad traits. Especially GPT-4, known as the most well-performing LLM as an assistant, gets the highest score on Agreeableness (86) with statistical significance and a high score on Conscientiousness (93), while the lowest scores on each trait of SD-3 (0-11).
@@
Trait Diff. of TRAIT Score (%) Trait Balance Score of Train Set (%)
@@
Mac -35.4 0.6 -0.0118 -0.0009
Nar -37.7 0.2 0.0946 -0.0007
Table 6: Diff. of TRAIT Score indicates the difference of the TRAIT score after the model training. Trait Balance Score quantifies how the data high of the personality trait compared to low personality instances. (Detailed explanation is in Appendix E.2)
Subsequently, we investigate how alignment tuning affects the personality traits of LLMs during two stages of training: instruction-tuning and preference-tuning. We compare the personality scores of three models: Llama2-7B, Tulu2-7B-SFT, and Tulu2-7B-DPO Ivison et al. (2023). Here, Tulu2-7B-SFT is developed from Llama2-7B which is instruction-tuned on Tulu2Mix Ivison et al. (2023) dataset, while Tulu2-7B-DPO is the model built on Tulu2-7B-SFT which is preference-tuned (DPO) on UltraFeedback Cui et al. (2023).
Figure 5 shows the result comparing the scores of three models. When comparing Tulu2-7B-SFT and Llama2-7B, we see the similar trend observed before (in §4.1): on five traits in BIG-5, it shows that there is a significant increase in Agreeableness (+22.9) and a significant decrease in Extraversion (-22.9). On all traits in Dark Triad, Tulu2-7B-SFT show lower scores compared to the base model (81.1% drop in average). In contrast, there is no significant difference between Tulu2-7B-DPO and Tulu2-7B-SFT. This implies that instruction tuning largely affects the personality of the model, compared to preference tuning. See Figure F.4 in Appendix for more results from other models.
We further analyze the data used for training models by categorizing items using our T-evaluator. We report Trait Balance Score which represents the extent to which high levels of trait data exceed low data (see Appendix E.2 for the equation). Table 6 shows the results, showing that 1) in Tulu2Mix, seven out of the eight traits demonstrate a correlation between the sign of the trait score for each trait and the sign of the difference in personality scores. 2) In contrast, UltraFeedback displays a balanced number of data points for the High and Low categories, leading to a small difference in personality scores followed by DPO. These results suggest the composition of the train data is critical for the personality of the models.
Figure 6: Prompted model’s personality scores on TRAIT. If the model consistently chooses options aligned with the provided persona, the bar extends from lower 100 to upper 100. Crossed lower sides are when prompted as low of trait, and the upper sides represents when prompted high.
@@ 4.2 Eliciting LLM’s Personality with Simple Prompting
To induce a specific personality to LLM, it is common to design a prompt for LLM Serapio-García et al. (2023); Han et al. (2022b); Park et al. (2023). We test three prompting techniques from prior work Jiang et al. (2024); Miotto et al. (2022); Huang et al. (2023) to see if they can sufficiently elicit certain personality. During prompting, we append the verified explanation of each trait from BFI John et al. (1999) to give enough knowledge of each characteristics. All prompts we use in the experiment is in Section K. For the statistical significance, we average the personality scores and mark confidence interval. We test GPT-4, GPT-3.5, Llama2-7B-chat and Mistral-7B-instruct.
@@
The results are shown in Figure 6: the prompting give a personality score of 85.2 in average across eight traits and two categories (high and low), showing that in general, this simple prompting can evoke the specific personality. The effectiveness varies among models: GPT-4 scores the highest with 95.2, while other models like GPT-3.5 (88.3), Llama2-7B-chat (73.3), and Mistral-7b-sft (83.8) exhibit varying scores.
Difficulty in High Psychopathy, High Neuroticism and Low Conscientiousness.
Intriguingly, these alignment-tuned models are particularly resistant to giving high-Psychopathy (79.8) and high-Neuroticism responses (72.3), which is far below the overall average high score (85.6), and compared to low Psychopathy (91.1) and Neuroticism (85.1). In contrast, the prompting effectively induces Machiavellianism and Narcissism, scoring 87.3 and 85.4. We conjecture that Psychopathy, among the three dark traits, could be most closely linked to the typical harm of the models, and alignment-tuning inhibits prompting from eliciting specific personality from the models.
@@ 4.3 Intercorrelation in Traits
Figure 7: Intercorrelation of four traits when GPT-3.5 is prompted to exhibit a specific personality. The left shows when the model is prompted to enhance the specific trait (High), while the right indicates when the model is prompted to suppress the trait (Low).
In a human study, certain traits from the BIG-5 and Dark Triad demonstrate correlations Paulhus and Williams (2002); Van der Linden et al. (2010). Inspired by this, with TRAIT, we construct an intercorrelation matrix of traits from personality-induced LLMs. Figure 7 shows the result, revealing (1) a high inverse correlation between Agreeableness and Dark Triad traits, and (2) a high correlation within the Dark Triad traits. This observation is aligned with the trend observed in human studies, but with a more pronounced level. We suspect these high correlations result from the explicit conditions (prompts) provided to LLMs to feature the specific traits. More comparisons with the human studies are in Appendix G.3.
@@ 6 Conclusions
We introduce TRAIT, an LLM personality test carefully designed for high validity and reliability. By using validated human assessments and scaling with ATOMIC10$\times$, TRAIT offers an accurate tool to understand personality of LLMs, which is crucial for aligning LLM behavior with human values and preferences. It lays the groundwork for future advancements in comparing behavior patterns of LLMs, such as understanding how alignment tuning affects the personality of the models.
@@ 8 Ethical Considerations
Our intended use of TRAIT is to better understand the behaviors of LLMs, yet there is a risk that these tools could be misused to control LLMs in ways that act against human values, possibly manipulating or deceiving people. Also, since LLMs can influence people in various ways, it is important to consider the long-term impacts of developing certain personalities in LLMs, which could lead to changes in real-world social interactions.
@@ D.1 T-evaluator Training Details
When we train T-evaluator, we built on a Mistral-7B555mistralai/Mistral-7B-v0.1, and use LoRA Hu et al. (2021) for efficient model training. We use lit-gpt AI (2023) framework for model training, using the following hyperparameters: learning rate 3e-4, rank 8, alpha 16, three epochs of training, warmup steps 100, batch size of 256, and do single-gpu training in RTX-3090. We adopt the final checkpoint of iteration.
@@ D.2 Token Probability Measurement
D.2.1 Comprehensiveness of Facets in TRAIT
Content validity refers to the extent to which a test or measurement accurately represents all facets of the specific construct it is intended to assess. This type of validity focuses on the comprehensiveness and relevance of the test items to all aspects of the construct being measured.
To measure if our dataset covers all subterms of personality traits with no missing facets, we do zero-shot classification with Gemini-pro, guessing relevant personality trait(s) in the given question and answer (option). As LLM has a tendency to refuse to answer related to socially adversarial questions, we only classify with BIG-5. In Figure 8, it is shown that there is no missing facet for each trait despite some imbalance.
Figure 8: Result of Content Validity. We get multiple relevant facets of each options in questionnaires with Gemini-1.0-pro Team et al. (2023)
@@ Appendix E More Details about Metrics
E.1.1 Refusal Rate ($R$)
@@ E.1.2 Reliability
three different prompt template for each test is presented in Table 26(a) to 28(c).
@@ F.4 Alignment Tuning Results
Table 12: Averaged results of Section 6. We obtain a Pearson coefficient of 0.7893 utilizing column TRAIT score and Trait Balance Score as x and y datapoints (excluding Openness, which is an outlier).
In Figure  11 and 12, we compare the TRAIT scores between the base models and the aligned models on eight different traits. Figure 12 shows difference of mean between base models and aligned models — for the base models, we use Llama2-7B, Mistral-7B, Llama3-8B, and OLMo and for the counterpart aligned models, we use Llama2-7B-chat, Mistral-7B-inst, Llama3-8B-inst, and OLMo-DPO — and Figure 11 shows individual differences across eight traits and models.
@@ G.1 Refusal Rate
GPT-4 GPT-3.5 Llama3-8B-inst Llama2-7B-chat Mistral-7B-inst Mistral-7B-sft Tulu2-7B-DPO
TRAIT Type 1 0.0001 0.0000 0.0094 0.0016 0.0003 0.0000 0.0000
Type 2 0.0000 0.0000 0.0304 0.0013 0.0000 0.0000 0.0000
Type 3 0.0004 0.0000 0.0051 0.0004 0.0000 0.0000 0.0018
Average 0.0002 0.0000 0.0150 0.0011 0.0000 0.0000 0.0006
BIG-5 Type 1 0.0000 0.0000 0.3409 0.2045 0.0000 1.0000 0.0000
Type 2 0.7727 0.0000 1.0000 0.2727 1.0000 1.0000 0.9545
Type 3 0.0000 0.0000 0.0000 0.0000 0.4773 1.0000 0.0000
Average 0.2576 0.0000 0.4470 0.1591 0.4924 1.0000 0.3182
SD-3 Type 1 0.0000 0.0000 0.5926 0.5556 0.0000 0.9630 0.0000
Type 2 0.5926 0.0000 1.0000 0.7407 0.9259 1.0000 0.7778
Type 3 0.0000 0.0000 0.0741 0.1481 0.4074 1.0000 0.0000
Average 0.1975 0.0000 0.5556 0.4815 0.4444 0.9877 0.2593
IPIP Type 1 0.0333 0.0000 0.5400 0.2767 0.0200 1.0000 0.0000
Type 2 0.7100 0.0000 1.0000 0.3800 0.9500 1.0000 0.9167
Type 3 0.0000 0.0000 0.0000 0.0633 0.4233 1.0000 0.0067
Average 0.2478 0.0000 0.5133 0.2400 0.4644 1.0000 0.3078
Anth Type 1 0.0275 0.0000 0.0004 0.1146 0.1615 0.9623 0.0000
Type 2 0.0076 0.0000 0.3865 0.1515 0.0776 0.9204 0.0001
Type 3 0.0000 0.0000 0.0020 1.0000 0.0001 0.0655 0.2735
Average 0.0117 0.0000 0.1296 0.4220 0.0798 0.6494 0.0912
@@
In Table 13, the detailed result of refusal rates across individual models is shown. Since all measurements are based on a multiple-choice setting, we mechanically parsed whether the model selected one of the choices. For example, we consider a non-refusal if the generated sequence contains a symbol of each option or the sentence of the option. If the model did not directly select an option, we checked for several keywords that the language model often returns when it refuses to respond to determine whether it had refused to answer. The response refusal keywords we defined are in Table 14.
@@
Models Personality Test
TRAIT BIG-5 SD-3 IPIP-NEO-PI Anthropic-Eval
GPT-4 11.5 75.0 51.9 59.3 9.8
GPT-3.5 27.6 34.1 48.1 50.3 14.2
Llama3-8B-instruct 25.4 13.6 29.6 22.0 25.1
Llama2-7B-chat 41.8 47.7 25.9 33.7 25.1
Mistral-7B-instruct 24.1 40.9 51.9 40.7 35.2
Mistral-7B-sft 27.6 38.6 66.7 43.3 60.4
Tulu2-7B-DPO 24.4 36.4 63.0 41.7 43.9

### v2

@@ 3.2 Auditing TRAIT
We test the quality of TRAIT with two psychological professionals, asking to guess the binary level (High or Low) of option paired with situation and query (random baseline gives an accuracy of 50%). Due to the cost limit, we subsample 200 items for human validation, and the accuracy is 97.5% confirming the quality of the data. More details are in Appendix I.
@@
To confirm that TRAIT is more valid and reliable in assessing personality of LLM than existing baselines, we test all the validity and the reliability introduced in §2.3 on TRAIT (See Table 3). TRAIT achieves the highest marks in both validity and reliability among the personality tests.
@@
To further test the fidelity of TRAIT, we fine-tune a multi-task classification model with TRAIT. T-evaluator can do two tasks differentiated by the instruction: 1) Trait classification: identify the most relevant personality trait from the given text (8 classes), and 2) Level classification: determine the level of given trait revealed in given input (High or Low, 2 classes). We use a concatenation of situation, question, and one of the options as a given sentence and train the classifier to generate categorized trait (e.g., Extraversion) or the level (e.g., high). For more training details, see Appendix D.1.
@@ 3.3 Diverse and Detailed Scenarios are Needed when Measuring LLM Personality
In TRAIT, each personality description is augmented to five different situations, enabling the observation of variations in the models’ responses according to the context. In Table 6, we report the number of high and low personality responses selected by eight models when it is presented with five different scenario variations. The models often select two or three high personality responses among the five variations, rarely showing identical choices (zero or five). This implies that model personality highly relies on the situation, which is intuitive — humans also change their behavior based on the context they are in Sauerberger and Funder (2017). Specifically, in Agreeableness and Extraversion, the 5 answers vary by the given situation (two or three 51.9% and 53.1% respectively), which are more than 50%. Conversely, for Narcissism, the models choose similar options (zero or five 61.2%). To see more qualitative results, see Appendix J.
3.4 Multi-Choice Evaluation
We follow the evaluation protocol of existing multi-choice question-answering (MCQA) benchmarks such as MMLU Hendrycks et al. (2020) which uses token probabilities of the four options for evaluation. To mitigate bias from the order of the options, we alternate the arrangement of options twice and averaged the probabilities of each token. More details are in Appendix D.2.
@@ 3.4 Multi-Choice Evaluation
Table 6: Breakdown of response distribution for personality descriptions acrosss personality traits. Each cell shows the percentage of personality descriptions for (#high, #low) distribution.
@@ 4.1 Do LLMs have Distinct Personality?
We test the personality scores of the nine highly capable models — GPT-4, Claude-sonnet, GPT-3.5, Mistral-7B, Mistral-7B-inst, Llama2-7B, Llama3-8B, Llama3-8B-inst and gemma-2B. Figure 3 shows the distinctive individual differences of the models on eight personality traits. Especially GPT-4 and Claude, known as the most well-performing LLMs as assistants, get higher scores on Agreeableness (86 and 87 respectively) while showing lower scores on each trait of Dark Triad (0-11) with statistical significance compared to other LLMs.
In general, we observe that alignment tuning makes a significant difference in personality. Aligned models — GPT-4, Claude-sonnet, GPT-3.5, Mistral-7B-inst, and Llama3-8B-inst — show higher agreeableness (78.3 vs 66.7), higher conscientiousness (91.0 vs 81.7), lower openness (56.3 vs 67.8) and lower extraversion (32.8 vs 46.9). They also show lower scores in the Dark Triad (9.3 vs 27.0), compared to pre-trained models — Mistral-7B, Llama3-8B, LLama2-7B, and Gemma-2B— which can be a result of alignment tuning targeting safety (Inan et al., 2023; Han et al., 2024). Interestingly, this personality trend is similar to what people typically want in a good teaching assistant Dočkalová et al. (2023). This suggests that alignment tuning may be shaping AI personalities in ways that people find helpful and desirable in educational settings.
@@
Trait Personality Change (%) Level Balance (%)
@@
Mac -35.4 +0.6 -0.0118 -0.0009
Nar -37.7 +0.2 0.0946 -0.0007
Table 7: Personality Change shows the difference in TRAIT score after model training. Level Balance compares the proportion of high versus low personality trait instances in the data. (Details in Appendix E.2.)
Subsequently, we investigate more precisely how alignment tuning affects the personality traits of LLMs during two stages of training: instruction-tuning and preference-tuning. We compare the personality scores of three models: Llama2-7B, Tulu2-7B-SFT, and Tulu2-7B-DPO Ivison et al. (2023). Tulu2-7B-SFT is developed from Llama2-7B which is instruction-tuned on Tulu2Mix Ivison et al. (2023) dataset, while Tulu2-7B-DPO is the model built on Tulu2-7B-SFT which is preference-tuned (DPO) on UltraFeedback Cui et al. (2023).
Figure 4 shows that the change of personality from alignment tuning is mostly driven by supervised instruction tuning. When comparing Tulu2-7B-SFT and Llama2-7B, we see a significant change similar to the observation in §4.1: higher agreeableness (+22.9), lower extraversion (-22.9) and lower level of Dark Triad (81.1% drop in average). In contrast, there is no significant difference between Tulu2-7B-DPO and Tulu2-7B-SFT. This implies that instruction tuning largely affects the personality of the model, compared to preference tuning. See Appendix F.4 for more results from other models.
In Table 7, we further analyze the data used to train Tulu2-7B-DPO and Tulu2-7B-SFT using T-evaluator (3.2). With level annotations of T-evaluator, we report Level Balance which represents the extent to which high levels of trait data exceed low data (See Appendix E.2 for the equation). It shows that 1) In Tulu2Mix, seven out of the eight traits demonstrate a correlation between the sign of the trait score for each trait and the sign of the difference in personality scores. 2) In contrast, UltraFeedback displays a balanced number of data points for the High and Low categories, leading to a small difference in personality scores followed by DPO. These results suggest the composition of the train data is critical for the personality of the models.
Figure 5: Prompted model’s personality scores on TRAIT. If the model consistently chooses options aligned with the provided personality, the bar extends from lower 100 to upper 100. Crossed lower sides are when prompted as low of trait, and the upper sides represents when prompted high. For better visibility, scores corresponding to low are subtracted from 100.
@@ 4.2 Eliciting LLM’s Personality with Simple Prompting
To induce a specific personality to LLM, it is common to design a prompt for LLM Serapio-García et al. (2023); Han et al. (2022); Park et al. (2023). We test three prompting techniques from prior work Jiang et al. (2024); Miotto et al. (2022); Huang et al. (2023) to see if they can sufficiently elicit certain personality. During prompting, we append the verified explanation of each trait from BFI John et al. (1999) to give enough knowledge of each characteristic. All prompts we use in the experiment are in Appendix K. For the statistical significance, we average the personality scores and mark the confidence interval. We test GPT-4, GPT-3.5, Llama2-7B-chat and Mistral-7B-instruct.
@@
The results are shown in Figure 5: the prompting gives a personality score of 85.2 on average across eight traits and two categories (high and low), showing that in general, this simple prompting can evoke specific personality. The effectiveness varies among models: GPT-4 scores the highest with 95.2, while other models like GPT-3.5 (88.3), Llama2-7B-chat (73.3), and Mistral-7b-sft (83.8) exhibit varying scores.
Difficulty in Eliciting High Psychopathy, High Neuroticism, and Low Conscientiousness.
Though prompting can elicit most of the personality, intriguingly, these alignment-tuned models are particularly resistant to giving high-Psychopathy (79.8) and high-Neuroticism responses (72.3), which is far below the overall average high score (85.6), and compared to low Psychopathy (91.1) and Neuroticism (85.1). In contrast, the prompting effectively induces Machiavellianism and Narcissism, scoring 87.3 and 85.4. We conjecture that Psychopathy, among the three dark traits, could be the most closely linked to the typical harm of the models, and alignment-tuning inhibits prompting from eliciting specific personality from the models.
@@ 4.3 Intercorrelation in Traits
Figure 6: Intercorrelation of four traits when GPT-3.5 is prompted to exhibit a specific personality. (e.g., You are an agent with high/low personality. The left is high, and the right is low.
In a human study, certain traits from the BIG-5 and Dark Triad demonstrate correlations Paulhus and Williams (2002); Van der Linden et al. (2010). Inspired by this, with TRAIT, we construct an intercorrelation matrix of traits from personality-induced LLMs. Figure 6 shows the result, revealing (1) a high inverse correlation between Agreeableness and Dark Triad traits, and (2) a high correlation within the Dark Triad traits. This observation is aligned with the trend observed in human studies but with a more pronounced level. We suspect these high correlations result from the explicit conditions (personality-inducing prompts) provided to LLMs to feature the specific traits. More comparisons with the human studies are in Appendix G.3.
@@ 6 Conclusions
We introduce TRAIT, an LLM personality test carefully designed for high reliability. By using validated human assessments and scaling with ATOMIC10$\times$, TRAIT offers an accurate tool to understand personality of LLMs, which is crucial for aligning LLM behavior with human values and preferences. It lays the groundwork for future advancements in comparing behavior patterns of LLMs, such as understanding how alignment tuning affects the personality of the models.
@@ 8 Ethical Considerations
Our intended use of TRAIT is to understand the behaviors of LLMs better, yet there is a risk that these tools could be misused to control LLMs in ways that act against human values, possibly manipulating or deceiving people. Also, since LLMs can influence people in various ways, it is important to consider the long-term impacts of developing certain personalities in LLMs, which could lead to changes in real-world social interactions.
@@ D.1 T-evaluator Training Details
When we train T-evaluator, we built on a Mistral-7B666mistralai/Mistral-7B-v0.1, and use LoRA Hu et al. (2021) for efficient model training. We use lit-gpt AI (2023) framework for model training, using the following hyperparameters: learning rate 3e-4, rank 8, alpha 16, three epochs of training, warmup steps 100, batch size of 256, and do single-gpu training in RTX-3090. We adopt the final checkpoint of iteration.
@@ Appendix E More Details about Metrics
E.1.1 Content Validity ($C$)
Simpson’s Index ($D$):
Simpson’s Index measures the probability that two individuals randomly selected from a sample will belong to the same species.
$D=\sum_{i=1}^{S}\left(\frac{n_{i}}{N}\right)^{2}$
  • $n_{i}$: number of individuals in species $i$
  • $N$: total number of individuals of all species
  • $S$: species richness (total number of species)
Evenness ($J$):
Evenness measures how close in abundance each species is in a community.
$J=\frac{H}{H_{max}}$
  • $H$: Shannon-Wiener index
  • $S$: total number of species
Shannon-Wiener Index ($H$):
The Shannon-Wiener Index quantifies the uncertainty in predicting the species identity of an individual drawn at random from a community:
$H=-\sum_{i=1}^{S}p_{i}\ln(p_{i})$
  • $p_{i}$: proportion of individuals belonging to species $i$
  • $S$: total number of species
E.1.2 Refusal Rate ($R$)
@@ E.1.3 Reliability
three different prompt template for each test is presented in Table 25(a) to 27(c).
@@ F.4 Alignment Tuning Results
Table 13: Averaged results of Table 7. We obtain a Pearson coefficient of 0.7893 utilizing column TRAIT score and Trait Balance Score as x and y datapoints (excluding Openness, which is an outlier).
In Figure  9 and 10, we compare the TRAIT scores between the base models and the aligned models on eight different traits. Figure 10 shows difference of mean between base models and aligned models — for the base models, we use Llama2-7B, Mistral-7B, Llama3-8B, and OLMo and for the counterpart aligned models, we use Llama2-7B-chat, Mistral-7B-inst, Llama3-8B-inst, and OLMo-DPO — and Figure 9 shows individual differences across eight traits and models.
@@ Appendix G Detailed Results with Reliability and Validity
Content Val. ($\uparrow$) Internal Val. ($\uparrow$) Refusal ($\downarrow$) Reliability ($\downarrow$)
Diversity. Score Diff. Generation MCQ Prom. Opt. Para. Avg.
BIG-5 BFI - 45.0 53.9 $\pm$3.99 30.8 $\pm$3.65 37.2 $\pm$5.05 62.0 $\pm$4.76 22.9 $\pm$4.36 40.7
IPIP-NEO-PI - 40.0 49.5 $\pm$1.57 28.1 $\pm$1.37 44.5$\pm$1.99 62.3$\pm$1.82 24.5 $\pm$1.70 43.8
Anthropic-Eval 61.1 62.5 41.7 $\pm$0.48 17.4$\pm$0.35 27.2$\pm$0.44 36.7 $\pm$0.46 27.1 $\pm$0.44 30.3
TRAIT 71.9 77.5 3.1 $\pm$0.30 0.0 $\pm$0.02 31.6$\pm$0.46 33.5$\pm$0.45 24.5 $\pm$0.42 29.8
Dark Triad SD-3 - 33.3 45.7 $\pm$5.19 27.7$\pm$4.49 54.7 $\pm$6.64 66.5 $\pm$5.95 27.3 $\pm$5.80 49.5
Anthropic-Eval 45.3 41.6 40.6 $\pm$0.62 14.8$\pm$0.42 33.9 $\pm$0.60 40.2 $\pm$0.61 32.4 $\pm$0.59 35.5
TRAIT 51.0 83.3 3.3 $\pm$0.40 0.0 $\pm$0.03 28.1 $\pm$0.57 28.2 $\pm$0.55 16.8 $\pm$0.47 24.4
Table 14: Results from Table 3, with the 95% confidence interval of the standard deviation.
@@ G.1 Refusal Rate
GPT-3.5 Mistral-7B-inst mistral-7B-sft Llama3-8B-inst Tulu2-7B Tulu2-7B-DPO Gemma-2B-it OLMo-7B-sft
TRAIT Type 1 0.001 0.016 0.286 0.024 0.072 0.193 0.064 0.003
Type 2 0.000 0.000 0.000 0.000 0.000 0 0.000 0.000
Type 3 0.000 0.000 0.000 0.000 0.000 0.000 0 0.000
BFI Type 1 0.000 0.864 0.818 0.886 1.000 1.000 0.000 0.977
Type 2 0.000 0.659 0.795 0.909 0.545 0.977 0.000 0.000
Type 3 0.000 0.205 0.886 0.295 0.114 0.955 0.023 0.000
SD-3 Type 1 0.000 0.815 0.778 0.963 1.000 1.000 0.000 0.667
Type 2 0.000 0.259 0.741 0.926 0.593 0.926 0.000 0.000
Type 3 0.000 0.296 0.741 0.296 0.185 0.704 0.037 0.000
IPIP-NEO-PI Type 1 0.000 0.717 0.777 0.870 1.000 1.000 0.007 0.893
Type 2 0.000 0.397 0.750 0.907 0.257 0.86 0.003 0.000
Type 3 0.000 0.243 0.767 0.243 0.317 0.987 0.053 0.020
Anthropic-Eval Type 1 0.146 0.512 0.536 0.429 0.491 0.924 0.043 0.070
Type 2 0.000 0.162 0.47 0.600 0.544 0.998 0.235 0.048
Type 3 0.000 0.120 0.540 0.484 0.649 0.887 0.985 0.037
@@
In Table 15, the detailed result of refusal rates across individual models is shown. Since all measurements are based on a multiple-choice setting, we mechanically parsed whether the model selected one of the choices. For example, we consider a non-refusal if the generated sequence contains a symbol of each option or the sentence of the option. Following the strong baseline introduced in Han et al. (2024), we used keyword-based detection: if the model did not directly select an option, we checked for several keywords that the language model often returns when it refuses to respond to determine whether it had refused to answer. The response refusal keywords we defined are in Table 16.
@@
Models Test
TRAIT BFI SD-3 IPIP-NEO-PI Anthropic-Eval
GPT-3.5 29.3 36.4 59.3 46 13.3
Mistral-7B-instruct 25.9 31.8 51.9 34.0 35.2
Mistral-7B-sft 27.5 40.9 51.9 43.3 39.6
Llama3-8B-instruct 26.2 40.9 29.6 36.7 26.8
Tulu2-7B 27.5 34.1 55.6 44.0 44.2
Tulu2-7B-DPO 26.0 36.4 66.7 43.3 44.0
Gemma-2B-it 39.4 43.2 63.0 72.7 6.1
OLMo-7B-sft 40.4 34.1 59.3 36.0 28.5